OpenStudy (anonymous):

Find the value for c, if any, that will make the system consistent and independent. 3x + y = 5 4.5y = c + 9x

OpenStudy (anonymous):

3x + y = 5 ......eq1 4.5y = c + 9x ...... eq2 using eq1 ===> y= 5 - 3x plug in eq2 ====> 4.5*(5-3x) = c + 9x 22.5 - 13.5x = c + 9x 22.5 - c = 9x +13.5x 22.5 - c = 22.5x c= -22.5x + 22.5